### v2

# A Note on Power-OTMs

Abstract
We consider the computational strength of Power-OTMs, i.e., ordinal Turing machines equipped with a power set operator, and study a notion of realizability based on these machines. When parameters are allowed, these machines are equivalent to the Set Register Machines defined by Robert Passmann in [14], and most of the results on Power-OTMs with parameters are analogous to results obtained by Passmann. New are the results about the strength of power-OTMs without parameters.
1 Introduction
Ordinal Turing Machines (OTMs) are a model of transfinite computability introduced by Peter Koepke in [10]. A notion of realizability based on OTMs was considered in [5]; the guiding idea here is to consider OTM-computable functions as the underlying notion of effectivity on which the realizability notion is then based. A basic result on OTMs is that (provably in ZFC), OTMs cannot compute power sets. However, it has been argued that the formation of power sets should be regarded as a basic constructive operation in set theory, and realizability notions based on this idea have been defined and studied, see, e.g., Rathjen [13], p. 3 or Moss [12], p. 248. This motivates the study of OTMs equipped with an extra command for forming power sets, which we call power-OTMs. After finishing this note, we became aware that a very similar investigation was conducted earlier by Robert Passmann in [14], whose “Set Register Machines” are essentially a register variant of power-OTMs, designed to avoid coding issues (the equivalence of the two models when the power-set operator is dropped is proved in [14], Theorem 3.8). Below, we will indicate the corresponding results about SRMs. It still seems to be worth noting that these results can be integrated in the more familiar OTM-setting. Passmann allows his SRMs the use of arbitrary set parameters; in contrast, we consider only the cases where (i) a single ordinal parameter and (ii) no parameters are allowed. The results concerning the latter seem to be essentially new. Moreover, unlike Passmann, we have no use for the assumption of a global well-ordering, because the encoding of sets as an OTM-input induces a well-ordering anyway.
2 Basic Definitions and Results
We assume that the reader is familiar with Koepke’s Ordinal Turing Machines (OTMs), see [10] or [4], section 2.5.6. For convenience, we will consider OTMs with four tapes: An input tape, a scratch tape, an output tape and an extra tape for carrying out the operations to which we are reducing.
Definition 1
Let $F:V\rightarrow V$ be a class function. An $F$-encoding is a map $\hat{F}:\mathfrak{P}(\text{On})\rightarrow\mathfrak{P}(\text{On})$ which, for every $c_{x}\subseteq\text{On}$ that codes a set $x$, maps $c_{x}$ to a code $c_{F(x)}\subseteq\text{On}$ for $F(x)$. An $F$-OTM-program $P$ is an OTM-program using four tapes that has as an additional command an $F$-command. We call the four tapes the input tape, the scratch tape, the output tape and the $F$-tape.
If $\hat{F}$ is an $F$-encoding, then the computation of $P$ (on a given input, $x$ if there is one) is obtained by running the program as usual (according to the definition of OTM-computations in [10]), and, whenever the $F$-command is carried out, replacing the current content $c$ of the $F$-tape by $\hat{F}(c)$. We denote this computation by $P^{\hat{F}}(x)$.
If, for all inputs, the set encoded by the output of $P$ is the same for all encodings of $F$ and all encodings of the input $x$, we call $F$ coding-stable. [^1] Similar notions are used in [5] (there called uniformity) or [1]. In this case, we will denote the set encoded by the output of $P^{\hat{F}}(c)$ for any encoding $\hat{F}$ of $F$ and any code $c_{x}$ for a set $x$ by $P^{F}(x)$.
A set $X$ is called $F$-OTM-computable if there are a coding-stable $F$-OTM-program $P$ (and perhaps an ordinal parameter $\rho\in\text{On}$) that halts with a code for $X$ on its output tape. $X$ is called $F$-OTM-recognizable [^2] The notion of recognizability was first defined by Hamkins and Lewis for Infinite Time Turing Machines in [9], see Theorem 4.9. if and only if there is a coding-stable OTM-program $P$ (and, perhaps, a parameter $\rho\in\text{On}$) such that, for all $s\subseteq\text{On}$:
$P^{F}(s,\rho)=\begin{cases}1\text{, if }s\text{ codes }x\\ 0\text{, otherwise}\end{cases}$
IF we can take $\rho=0$, we speak of parameter-free $F$-OTM-computability and -recognizability, respectively.
The definition of $F$-realizability is obtained from the definition of uniform OTM-realizability in [5] by replacing OTMs with set parameters by coding-stable $F$-OTMs everyhwere.
Similarly, the definition of recognizable $F$-realizability, or $F$-$r$-realizability, is obtained from the definition of $r$-realizability in [3] by replacing OTMs with coding-stable $F$-OTMs everywhere.
If $F$ is the power set operator $x\mapsto\mathfrak{P}(x)$, we speak of power-OTMs and power-realizability.
We fix a natural enumeration $(P_{i}:i\in\omega)$ of power-OTM-programs.
Remark 1
Power-OTM-computability is conceptually closely connected to Moss’ concept of “power-set recursion”, defined in [12]. We will see below that power-set recursion is strictly weaker than parameter-free power-OTM-computability. Moreover, Moss does not consider realizability in his work, which is the main motivation of the present paper.
An important feature of OTMs is their ability to enumerate Gödel’s constructible universe $L$. Similarly, Power-OTMs can enumerate the set-theoretical universe $V$:
Lemma 1
[Cf. Passmann, [14]]
  1. 1.
    There is a (parameter-free) non-halting power-OTM-program $P_{\text{level-enum}}$ such that, for every ordinal $\alpha$, there is an ordinal $\tau$ such that, at time $\tau$, $P_{\text{level-enum}}$ has a code for $V_{\alpha}$ on its output tape. Moreover, there is a non-halting power-OTM $P_{\text{set-enum}}$ such that, for every set $x$, there is an ordinal $\tau$ such that, at time $\tau$, $P_{\text{set-enum}}$ has a code for $x$ on its output tape. [^3] This is essentially Passmann, [14], Proposition 3.12.
  2. 2.
    There is a (parameter-free) power-OTM-program $P_{\text{level}}$ such that, for every code $c_{x}$ for a set $x$, $P_{\text{level}}(x)$ halts with a code for the $\in$-minimal $V_{\alpha}\ni x$ on its output tape. [^4] See Passmann, [14], Proposition 3.10.
  3. 3.
    There is a (parameter-free) power-OTM-program $P_{\text{card}}$ which, given a code $c_{x}$ for a set $x$, computes $\text{card}(x)$. [^5] This is essentially the reverse direction of Passmann, [14], Proposition 3.10.
Proof
We give a brief sketch. (The arguments work as those of Passmann quoted above.) $P_{\text{level-enum}}$ is obtained by starting with the empty set, iteratively applying the power set operator, storing the intermediate results on the working tape and computing unions at limit times. $P_{\text{set-enum}}$ is then obtained from $L_{\text{level-enum}}$ by additionally running through each newly written level and writing codes for all of its elements to the output tape. If a set $x$ is given, one can search each level occuring in the run of $P_{\text{level-enum}}$ to check whether it contains $x$ and halt once the answer is positive, which yields $P_{\text{level}}$. Finally, for $P_{\text{card}}$, first use $P_{\text{level-enum}}$ to find some $V_{\alpha}\ni x$ that contains a bijection between $x$ and some ordinal for some given set $x$, and then search through $V_{\alpha}$ for the minimal ordinal for which there is such a bijection.
The ability of power-OTMs to enumerate the whole set-theoretical universe has the consequence that computability and recognizability coincide:
Theorem 1
A set is power-OTM-computable if and only if it is power-OTM-recognizable. In particular, power-realizability coincides with power-r-realizability.
Proof
As usual, computable sets can be recognized by computing (a code for) them and then comparing that to the oracle.
In the other direction, if $Q$ is a power-OTM-program (possibly with ordinal parameters) that recognizes a set $x$, we can compute $x$ by running $P_{\text{level-enum}}$ and checking all elements of all occuring levels with $Q$. Eventually, a code for $x$ will be produced and recognized, in which case the program writes that code to the output tape and halts.
This shows that the whole set-theoretical universe is “recursively enumerable” in the sense of Power-OTMs. Concerning the computational strength with ordinal parameters, we obtain the following:
Corollary 1
[Cf. Passmann, [14], Proposition 3.12] Every set is computable by Power-OTMs with ordinal parameters.
Without ordinal parameters, we need to work a bit more.
Definition 2
Let $\mathcal{L}_{\in,\mathcal{P}}$ denote the extension of the language of set theory with the one-place function symbol $\mathcal{P}$ (to be interpreted as the power set operator). [^6] This language was considered, e.g., by Rathjen, [13]. Let $\Sigma_{1}^{\mathcal{P}}$ denote the set of $\Sigma_{1}$-formulas in that language. [^7] Cf. ibid. For every $\Sigma_{1}^{\mathcal{P}}$-formula $\varphi$, define
$\alpha(\varphi):=\begin{cases}\text{the smallest ordinal }\alpha\text{ such that }V_{\alpha}\models\varphi\text{, if it exists}\\ 0\text{, otherwise.}\end{cases}$
Finally, let $\sigma^{\mathcal{P}}:=\text{sup}\{\alpha(\varphi):\varphi\in\Sigma_{1}^{\mathcal{P}}\}$. We fix a natural enumeration $(\psi_{i}:i\in\omega)$ of the $\Sigma_{1}^{\mathcal{P}}$-formulas.
Lemma 2
The set $\Phi:=\{\alpha(\varphi):\varphi\in\Sigma_{1}^{\mathcal{P}}\}$ has no maximum.
The following lemma is part of the folklore. We include a proof for the sake of completeness.
Lemma 3
Let $y$ be a set, $\psi$ be $\Delta_{0}$, $t:=\text{tc}(\{y\})$ and $\kappa:=\text{card}(t)$. If $\exists{x}\psi(x,y)$ then there exists some $x$ with $x\in V_{\kappa^{+}}$ so that $\psi(x,y)$.
Proof
Without loss of generality, let $\kappa\geq\omega$. By [11], Lemma 6.2, p. 131, this implies $t\in V_{\kappa^{+}}$. Now let $\beta$ be minimal such that $x\in V_{\beta}$; without loss of generality, we assume that $\beta>\kappa$. Let $H$ bet the transitive collapse of the $\Sigma_{1}$-hull of $t\cup\{x\}$ in $V_{\beta}$, and let $\pi:\Sigma_{1}\{t\cup\{x\}\}\rightarrow H$ be the collapse map. Then $\kappa\leq\text{card}(H)\leq\kappa\omega=\kappa$, so $\text{card}(H)=\kappa$. Moreover, $H$ is transitive, so we have $H\in H_{\kappa^{+}}$. [^8] Here, $H_{\kappa}$ refers, as usual, to the set of sets $x$ whose transitive closure has cardinality strictly less than $\kappa$. Since $t\subseteq H$, we have $\pi(y)=y$. Now $H\models\psi(\pi(x),y)$, and since $\psi$ is a $\Delta_{0}$-formula, we have $\psi(\pi(x),y)$ in $V$. Because of $\pi(x)\in H\in H_{\kappa^{+}}\subseteq V_{\kappa^{+}}$ (again because of [11], Lemma 6.2) and because $V_{\kappa^{+}}$ is transitive, we have $\pi(x)\in V_{\kappa^{+}}$.
Theorem 2
  1. 1.
    $\sigma^{\mathcal{P}}$ is the smallest ordinal $\alpha$ such that $V_{\alpha}$ contains all sets that are computable by a parameter-free power-OTM.
  2. 2.
    In particular, there are cofinally in $\sigma^{\mathcal{P}}$ many ordinals $\alpha$ such that $V_{\alpha}$ is power-OTM-computable without parameters.
  3. 3.
    $\sigma^{\mathcal{P}}$ is a limit cardinal with cofinality $\omega$.
  4. 4.
    There are a parameter-freely power-OTM-computable set $X$ and an $\in$-formula $\varphi$ such that $\{x\in X:\varphi(x,\rho)\}$ is not power-OTM-computable.
  5. 5.
    The set of power-OTM-computable sets is not transitive.
  6. 6.
    $V_{\sigma^{\mathcal{P}}}$ is admissible, but not $\Sigma_{2}$-admissible.
Proof
  1. 1.
    Let $P$ be a halting power-OTM-program. The proposition that $P$ halts is $\Sigma_{1}^{\mathcal{P}}$ and, by assumption, it holds in $V$; thus, by definition of $\sigma^{\mathcal{P}}$, it also holds in $V_{\sigma^{\mathcal{P}}}$. Thus, the whole computation of $P$ exists in $V_{\sigma^{\mathcal{P}}}$, and since $V_{\sigma^{\mathcal{P}}}$ is transitive, so does its output. Consequently, every power-OTM-computable set is an element of $V_{\sigma^{\mathcal{P}}}$.
    Conversely, let $\alpha<\sigma^{\mathcal{P}}$. By definition of $\sigma_{1}^{\mathcal{P}}$, there exists some $\Sigma_{1}^{\mathcal{P}}$-formula $\varphi$ such that $\alpha(\varphi)>\alpha$. But then, $V_{\alpha(\varphi)}$ is power-OTM-computable without parameters: To see this, use $P_{\text{level-enum}}$ and check, for every generated $V$-stage $V_{\beta}$, whether $V_{\beta}\models\varphi$. As soon as this happens to be true (which is guaranteed to eventually be the case), the program halts and outputs a code for $V_{\beta}$.
  2. 2.
    This follows immediately from the proof for (1).
  3. 3.
    Assume for a contradiction that $\sigma^{\mathcal{P}}$ is not a cardinal. Thus, there exists an ordinal $\alpha<\sigma^{\mathcal{P}}$ with $\text{card}(\alpha)=\text{card}(\sigma^{\mathcal{P}})$. By (2), there exists a (parameter-freely) power-OTM-computable stage $V_{\beta}$ with $\beta\geq\alpha$. Now $\text{card}(V_{\beta})\geq\alpha$. But if $V_{\beta}$ is power-OTM-computable without parameter, then so is $V_{\beta+1}=\mathcal{P}(V_{\beta})$, and thus, by Lemma 1, so is $\kappa:=\text{card}(V_{\beta+1})$. It follows that $\kappa\in V_{\sigma^{\mathcal{P}}}$, so $\kappa<\sigma^{\mathcal{P}}$, and this implies $\text{card}(\alpha)<\kappa\leq\sigma^{\mathcal{P}}$, a contradiction.
    Thus $\sigma^{P}$ is a cardinal. Concerning its cofinality, the function $f:\omega\rightarrow V_{\sigma^{\mathcal{P}}}$ with $f(k)=\alpha(\psi)$ is $\Sigma_{2}$ over $V_{\sigma^{\mathcal{P}}}$ and we have $\omega\in V_{\omega^{\mathcal{P}}}$. But, by definition of $\sigma^{\mathcal{P}}$, $f[\omega]$ is is unbounded in $\sigma^{\mathcal{P}}$. Hence $\text{cf}(\sigma^{\mathcal{P}})=\omega$.
  4. 4.
    Since the power set operator is definable by an $\in$-formula, the statement $k\in\omega\wedge\mathcal{P}_{k}\downarrow$ is expressable as an $\in$-formula $\varphi(k)$. If one uses separation with this formula from $\omega$, one obtains the halting set $h_{\mathcal{P}}$ for power-OTMs, which is not power-OTM-computable (which can be seen by the usual diagonal argument).
  5. 5.
    Clearly, $\mathcal{P}(\omega)$ is power-OTM-computable, but, by (4), the same does not hold for $h_{\mathcal{P}}\in\mathcal{P}(\omega)$.
  6. 6.
    Separation, Pot and AC hold in every $V$-stage $V_{\delta}$, provided $\delta$ is a limit cardinal; the same holds for the KP-axioms, with the exception of $\Delta_{0}$-Collection.
    To see $\Delta_{0}$-collection, consider a $\Delta_{0}$-formula $\psi(x,y)$ and a set $X\in V_{\sigma^{\mathcal{P}}}$ with $V_{\sigma^{\mathcal{P}}}\models\forall{x\in X}\exists{y}\psi(x,y)$. Pick $\gamma<\sigma^{\mathcal{P}}$ such that $x\in V_{\gamma}$ and $V_{\gamma}$ has a parameter-freely power-OTM-computable code (which exists by definition of $\sigma^{\mathcal{P}}$). Now $\text{card}(\text{tc}(x))\leq\text{card}(V_{\gamma})$ and $\delta:=\text{card}(V_{\gamma})$ is parameter-freely power-OTM-computable using $P_{\text{card}}$. It follows that $\delta^{+}$ is also parameter-freely power-OTM-computable, and hence, so is (a code for) $V_{\delta^{+}}$; so $V_{\delta^{+}}\in V_{\sigma^{\mathcal{P}}}$. But now, by Lemma 3, if, for some $x\in X$, there exists $y$ such that $\psi(x,y)$, then such $y$ exists in $V_{\delta^{+}}$. Thus $V_{\delta^{+}}$ is as required by $\Delta_{0}$-collection.
    In the proof for (3), we gave an example for a function $f:\omega\rightarrow V_{\sigma^{\mathcal{P}}}$ that is $\Sigma_{2}$-definable over $V_{\sigma^{\mathcal{P}}}$ but has $f[\omega]\notin V_{\sigma^{\mathcal{P}}}$.
We can now show that our notion of “computability with a power-set operator” is different from the one proposed by Moss [12]. Namely, in [12], power-recursive function are defined over so-called power-admissible sets, which are sets satisfying ZF with the replacement axiom restricted to $\Sigma_{1}$-formulas (see [12], p. 256).
Corollary 2
Parameter-free power-OTM-computability is strictly stronger than power-recursion.
Proof
The statement “There is a transitive model of power-KP” is clearly $\Sigma_{1}^{\mathcal{P}}$. Thus, the first power-admissible ordinal $\alpha$ is smaller than $\sigma^{\mathcal{P}}$.
In Moss [12] p. 260, it is shown that, provided strongly inaccessible cardinals exist, the first ordinal $\alpha$ such that $V_{\alpha}$ is closed under power-recursive functions is smaller than the smallest such. The same is not true for power-OTMs:
Lemma 4
Under the assumption that the respective cardinals exist, we have the following:
  1. 1.
    For every $n\in\omega$, $\sigma^{\mathcal{P}}$ is greater than the $n$-th strongly inaccessible cardinal.
  2. 2.
    $\sigma^{\mathcal{P}}$ is greater than the first strongly inaccessible limit of strongly inaccessible cardinals (i.e., the first $1$-inaccessible cardinal).
  3. 3.
    For every $k\in\omega$, $\sigma^{\mathcal{P}}$ is greater than the first $k$-inaccessible cardinal.
Proof
  1. 1.
    We enumerate $V$. For every ordinal $\alpha$ thus produced, we check (i) whether it is a regular cardinal (for this, it suffices to search through $\mathfrak{P}(\alpha)$) and (ii) whether $|\mathfrak{P}(\beta)|<\alpha$ for all $\beta<\alpha$ (which can be done by producing the power set of $\beta$ and then applying $P_{\text{card}}$ to it). By assumption, the answer will be positive at some point, at which we return $\alpha$, thus computing the first strongly inaccessible cardinal.
  2. 2.
    Using the routine described in the proof for (1), we can test a cardinal for being strongly inaccessible. We can then use this to search for a cardinal which is itself strongly inaccessible and has unboundedly many such cardinals below.
  3. 3.
    Inductively, there is a power-OTM-program for deciding the class of $k$-inaccessible cardinals, for every $k\in\omega$ (in fact, using recursion there is such a program that does this uniformly in $k$). Then again, we can search for the first cardinal that has this property.
Remark 2
The above could, of course, be further iterated.
3 Realizability with Power-OTMs
Realizability is defined by replacing OTMs with arbitrary set parameters with power-OTMs; we obtain two different notions of realizability, depending on whether ordinal parameters are allowed or not; however, all of the below results hold true for both of these notions, for much the same reasons. We therefore do not emphasize the distinction very much. Most of the results are, again, analogous to those obtained by Passmann in his investigation of SRM-realizability ([14], pp. 320–323).
Lemma 5
  1. 1.
    All axioms of KP are power-OTM-realizable. [^9] See Passmann, [14], Theorem 4.4.
  2. 2.
    The power set axiom is power-realizable. [^10] Passmann, [14], Theorem 4.4.
  3. 3.
    All instances of the $\Sigma_{1}$-separation scheme are power-realizable.
  4. 4.
    There are instances of the $\Sigma_{2}$-separation scheme which are not power-realizable. [^11] Cf. Passmann, [14], Theorem 4.5.
  5. 5.
    The set of power-OTM-realizable statements is closed under intuitionistic predicate logic, but not under classical predicate logic. [^12] Cf. Passmann, [14], Theorem 4.2.
Proof
  1. 1.
    The realizability of the KP axioms was already proved in [5] for ordinary OTMs; the arguments carry over verbatim to power-OTMs.
  2. 2.
    Trivial.
  3. 3.
    Let a $\Sigma_{1}$-formula $\varphi(y):\Leftrightarrow\exists{x}\psi(x,y)$ be given (where $\psi$ only contains bounded quantifiers), along with a set $X$. We want to compute (a code for) the set $\{y\in X:\varphi(y)\}$.
    The algorithm now works as follows: Let a set $y$ be given. We first compute the transitive hull $t:=\text{tc}(\{y\})$ (which is possible on a parameter-free standard OTM, i.e., without the power set operator), and, using Lemma 1(3), $\kappa:=\text{card}(t)$. By iterating the power set operator along $\kappa$, we compute a code for $V_{\kappa}$, and from this a code for $V_{\kappa^{+}}$. By Lemma 3, $V_{\kappa^{+}}$ contains some $x$ with $\psi(x,y)$, if one exists. Thus, it suffices to search $V_{\kappa^{+}}$ for such an $x$ in order to determine the truth value of $\exists{x}\psi(x,y)$. If so, $y$ becomes part of our set, otherwise, it does not.
  4. 4.
    As Passmann does it in his proof of [14], Theorem 4.5., we can use the formula expressing the $k$-th power-OTM-program will halt in the parameter $\rho\in\text{On}$. It only needs to be checked that this formula is in fact $\Sigma_{2}$, which is easily done.
  5. 5.
    The proof for OTM-realizability in [5] carries over without major modifications. Taking $\varphi(P,\rho)$ to be the statement “The power-OTM-program $P$, when run in the parameter $\rho$, will eventually halt or not halt”, it is clear that $\varphi(p,\rho)$ is classically a logical tautology, but not power-OTM-realizable (since a realizer would amount to a halting problem solver for power-OTMs on a power-OTM).
Remark 3
The same results hold verbatim for power-$r$-realizability, which, as we have seen, is identical to power-OTM-realizability.
4 Conclusion and further work
We have seen that regarding the power set operation as a “basic operation performed on sets” (Rathjen [13], p. 3) does indeed strongly increase the computational power of OTMs, but does not considerably alter their ability to realize impredicative instances of the separation scheme. It would be natural to consider notions of effective reducibility such as those considered in [1], [2] for OTMs based on power-OTMs. It would also be interesting to study the structure of the parameter-freely power-OTM-clockable ordinals. Finally, one could, in analogy with the work of Moss [12], study the structure of the parameter-free power-OTM-degrees.